OpenStudy (anonymous):

for x=2 and y=2 evaluate log(2)^2 and (log 2)^2 rounding to 3 decimal places

OpenStudy (rulnick):

log of 2^2 = log 4 = 0.602 (log 2)^2 = 0.091 ... assuming log is log base 10.